OpenStudy (anonymous):

What is the solution set of the absolute value sentence |3k + 4| = 8? A. {4/3, -4} B. {x|x > 4 or x < -4/5} C. No Solution D. {x|-11 < x < -7} E. All Real Numbers

OpenStudy (anonymous):

answer is A. {4/3, -4}

OpenStudy (anonymous):

(3k+4)= 8 3k= 8-4 k= 4/3 -(3k+4)=8 -3k-4=8 -3k = 8+4 k=-12/3 = -4 iesolution is -4, 4/3